OpenStudy (anonymous):

Store A is selling a video for 20% off the store’s regular price of $25.95. Store B is selling the same video for 30% off the store’s regular price of $29.50. Which store’s sale price is lower AND by how much lower is the price? a.Store A; $3.66 lower c.Store B; $3.66 lower b.Store A; $0.11 lower d.Store B; $0.11 lower

OpenStudy (anonymous):

@IQ250

OpenStudy (anonymous):

i need help! quick! @IQ250

OpenStudy (anonymous):

@Gabylovesyou

OpenStudy (anonymous):

@Adam96

OpenStudy (anonymous):

@ManWithQuestion

OpenStudy (gabylovesyou):

ok so first you should change the percents into decimals.. 20 / 100 = .2 30 / 100 = .3 subtract with 1.00 1.00 - .2 = .8 1.00 - .3 = .7 multiply each decimal with the regular prices 25.95 * .8 = 20.76 29.50 * .7 = 20.65 subtract each amount to get difference 20.76 - 20.65 = .11 Store B is lower with $.11 difference :)